I need this info, what is the  ORI, total asset turn over,  fixed assets, times interest earned ratio, debt ratio,  current ratio, average collection period, return on common equity,

inventory turnover ratio,  total asset turn over.

Jones CompanyBalance Sheet

  For the Year Ended 12/31/2015

Assets:

Cash and marketable securities $ 400,000

Accounts receivable 1,025,000

Inventories 1,937,500

Prepaid expenses   124,000

Total current assets $3,486,500

Fixed assets 2,800,000

  Less: accum. depreciation (1,087,500)

Net fixed assets $1,712,500

Total assets $5,199,000

Liabilities:

Accounts payable $ 340,000

Notes payable 825,000

Accrued taxes    42,000

Total current liabilities $1,207,000

Long-term debt 1,175,000

Owner's equity 2,817,000

Total liabilities and owner's equity $5,199,000

Jones Company Income Statement

For the Year Ended 12/31/2015


Net sales (all credit) $7,375,000

     Less: Cost of goods sold (4,312,500)

Selling and administrative expense (1,387,500)

Depreciation expense (135,000)

Interest expense (127,000)

Earnings before taxes 1,413,000

Income taxes (625,000)

Net income $ 788,000

Common stock dividends $488,000
Answer #1

Solution :-

(i) Total Assets Turnover Ratio = Net Sales / Total Assets = $7375000 / $5199000 = 1.418 times

(ii) Fixed Assets Turnover Ratio = Net Sales / Fixed Assets = $7375000 / $1712500 = 4.306 times

(iii) Times Interest Earned Ratio = EBIT / Interest Expense = $1540000 / $127000 = 12.126 times

EBIT = Earning before taxes + Interest = $1413000 + $127000 = $1540000

(iv) Debt ratio = Debt / (debt + Equity) = $1175000 / ( $1175000 + $2817000) = 0.294 : 1

(V) Current Ratio = Current Assets / Current Liabilities = $3486500 / $1207000 = 2.888 times

(vi) Accounts Receivable turnover Ratio = Net Credit Sales / Account Receivables = $7375000 / $1025000 = 7.195

Average Collection Period = 365 / Acc. Rec. turnover ratio = 365 / 7.195 = 50.72 Days

(vii) Return on Equity = Net Income / Equity = $788000 / $2817000 = 27.97%

(viii) Inventory Turnover ratio = Net Sales / Inventory = $7375000 / $1937500 = 3.806 times